OpenStudy (anonymous):

Two of the top grossing concert tours were by a jazz band and a rock band. Together the two tours visited 183 cities. The jazz band had visited 99 cities more than the rock band. How many cities did each group visit.

OpenStudy (kropot72):

Let the number of cities visited by the rock band = x Then the number of cities visited by the jazz band = x + 99 The total number of cities visited in the two tours were 183, therefore we can write the following equation: 183 = x + x + 99 .........(1) Now you just need to solve equation (1) to find x.

OpenStudy (anonymous):

Let j and r be the number of tour visits for the jazz and rock bands respectively. Solve the following for j and r: { j + r = 183, j - 99 = r } j = 141, r = 42